OpenStudy (anonymous):

A piece of art is in the shape of an equilateral triangle with sides of 6 in. Find the area of the piece of art. Round your answer to the nearest tenth.? A. 12.7 in.2 B. 15.6 in.2 C. 31.2 in.2 D. none of these

OpenStudy (anonymous):

anyone know this?

OpenStudy (anonymous):

@eSpeX do you know the answer to this?

OpenStudy (espex):

I believe the area of a triangle is \[\frac{1}{2} base \times height\]

OpenStudy (anonymous):

so 3 times 6 ????

OpenStudy (espex):

You don't provide a height in your problem so I assume you need to calculate it. My first instinct is to use trig but you would not have gotten to that yet. Does your class present you with SOH CAH TOA or any of the trig functions for finding a leg of a triangle?

OpenStudy (anonymous):

this is all online and no we work alone our teacher is a computer :( and i thought since all sides were six you just cut one six in half and times it by six?

OpenStudy (espex):

How about this: \[Area = \frac{\sqrt{3}}{4} \times side^2\]

OpenStudy (anonymous):

?????? what is that lol!

OpenStudy (espex):

If you measured the height you would have a new "leg" that was less than the 6, which was now a hypotenuse.